Oriole Company reported pretax net income from continuing operations of $1,000,000 and taxable
income of $1,200,000. The unfavorable book-tax difference of $200,000 was due to a $200,000
favorable temporary difference relating to depreciation, an unfavorable temporary difference of
$300,000 due to an increase in the reserve for bad debts, and a $100,000 unfavorable permanent
difference from the disallowance of compensation expense related to the exercise of incentive stock
options. Rate of tax is 21%

a. Compute current income tax expense

b. Compute deferred income tax expense or (tax benefit)

c.Compute the effective tax rate

d. Provide a reconciliation of hypothetical tax in dollars with its total tax provision in dollars
Hypothetical tax in dollars

e-Plus or minus difference in dollars

f-Equals total tax provision in dollars
Answer #1

Answer:

Requirements (a):

Pretax net income 1000,000
Favorable temporary difference (200,000)
Unfavorable temporary difference 300,000
Unfavorable permanent difference 100,000
Taxable income 12,00,000
Tax rate 21%
Current income tax expense 252000

Requirements-2

Favorable temporary difference (200,000)
Unfavorable temporary difference 300,000
Net unfavorable temporary difference 100,000
Tax rate 21%
Deferred tax assets 21000

Requirements-3

Total income tax provisions =252000-21000 = 231000

Effective tax rate =231,000/1000,000= 23.1%

Requirements-4

ETR reconciliation

Income tax expense at (21%)- hypothetical (1,000,000*21%) 210000
Tax cost from non deductible compensation (100,000*21%) 21000
Income tax provisions 231000
Hypothetical income tax rate 21.0%
Tax benefits from permanent difference 2.1%
Effective tax rate 23.1%